# Three-dimensional fracton topological orders with boundary Toeplitz braiding

### v4

@@ A.1 $\det B$=0
where $c_{2}\neq 0$ is a normalization constant. In this condition, only if $n=l_{1}=l_{2}=l_{4}=0$, $|m|<|l_{3}|$, there is a boundary zero mode $\mathbf{u}^{2}=\mathbf{w}^{2}$ given by Eq. (117), satisfying the boundary condition Eq. (95), where
$w^{2,a}=\frac{1}{\beta_{2}}\sqrt{\frac{1-\beta_{2}^{2N}}{1-\beta_{2}^{2}}},\quad w^{2,b}=0,\quad\beta_{2}=-\frac{l_{3}}{m}.$
In summary, if
$\left\{\begin{array}[]{l}m\neq 0,\quad n\neq 0,\\ l_{1}l_{4}=l_{2}l_{3},\\ 2l_{3}m=l_{4}n\ \text{or}\ 2l_{2}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\end{array}\right.$ (138)
@@
the $K$-matrix Eq. (2) has one boundary zero mode. More specifically, if $2l_{3}m=l_{4}n$, the boundary zero mode is a lower boundary zero mode, satisfying the boundary condition Eq. (94). If $2l_{2}m=l_{4}n$, the boundary zero mode is a upper boundary zero mode, satisfying the boundary condition Eq. (95). If
$\left\{\begin{array}[]{l}m\neq 0,\\ n=0,\\ l_{1}=l_{3}=l_{4}=0\ \text{or}\ l_{1}=l_{2}=l_{4}=0,\\ |m|<|l_{3}|\ \text{or}\ |m|<|l_{2}|,\end{array}\right.$ (143)
@@
the $K$-matrix has two boundary zero modes.
A.2 $\det B\neq 0$
If $\det B=l_{1}l_{4}-l_{2}l_{3}\neq 0$, Eq. (88) is a fourth-degree equation. If $\det B=\alpha_{1}=l_{1}l_{4}-l_{2}l_{3}\neq 0$, the solution to $\beta$ is
@@ A.2 $\det B\neq 0$
The solution gives
@@
If $l_{4}\neq 0$, there is a significant challenge in seeking a comprehensive analytical solution, to which a case-by-case study is necessary. To simplify discussion, we focus on the special case where $l_{4}=0$. Set $l_{4}=0$. Eq. (88) gives
$(l_{2}+m\beta+l_{3}\beta^{2})(l_{3}+m\beta+l_{2}\beta^{2})=0$ (145)
Since $\det B=0$ has been discussed, we assume that $\det B=l_{2}l_{3}\neq 0$. Solutions to Eq. (145) are
@@
Case I. $|\beta_{1}|>1$, $|\beta_{2}|<1$, $|\beta_{3}|>1$, $|\beta_{4}|<1$.
Case II. $|\beta_{1}|<1$, $|\beta_{2}|>1$, $|\beta_{3}|<1$, $|\beta_{4}|>1$.
Case III. $|\beta_{1}|>1$, $|\beta_{2}|<1$, $|\beta_{3}|<1$, $|\beta_{4}|>1$.
@@
The boundary bases corresponding to $\beta_{2}$ and $\beta_{4}$ are denoted as $\mathbf{w}^{2}$ and $\mathbf{w}^{4}$, respectively.
$\mathbf{w}^{2}=\begin{pmatrix}\beta_{2}w^{2,a}&\beta_{2}w^{2,b}&\beta_{2}^{2}w^{2,a}&\beta_{2}^{2}w^{2,b}&\cdots&\beta_{2}^{N}w^{2,a}&\beta_{2}^{N}w^{2,b}\end{pmatrix}^{T},$
$\mathbf{w}^{4}=\begin{pmatrix}\beta_{4}w^{4,a}&\beta_{4}w^{4,b}&\beta_{4}^{2}w^{4,a}&\beta_{4}^{2}w^{4,b}&\cdots&\beta_{4}^{N}w^{4,a}&\beta_{4}^{N}w^{4,b}\end{pmatrix}^{T}.$
@@
A trial solution $\mathbf{u}^{2,4}$
$\mathbf{u}^{2,4}=c_{2}\mathbf{w}^{2}+c_{4}\mathbf{w}^{4}=\begin{pmatrix}u^{2,4,A}_{1}&u^{2,4,B}_{1}&u^{2,4,A}_{2}&u^{2,4,B}_{2}&\cdots&u^{2,4,A}_{N}&u^{2,4,B}_{N}\end{pmatrix}^{T}$
is a linear combination of $\mathbf{w}^{2}$ and $\mathbf{w}^{4}$.
@@
We discuss the condition $\beta_{2}\neq\beta_{4}$ at first, which implies $m^{2}\neq 4l_{1}l_{2}$. Since $|\beta_{2}|<1$, $|\beta_{4}|<1$, $\mathbf{u}^{2,4}$ is required to satisfy the boundary condition Eq. (94), which renders
$\left(\begin{array}[]{cc}r_{1}&r_{2}\\ \beta_{2}^{2}l_{2}\left(l_{2}+\beta_{2}^{2}l_{3}+\beta_{2}m\right)+\beta_{2}m\left(l_{2}+\beta_{2}^{2}l_{3}+\beta_{2}m\right)&\beta_{4}^{2}l_{2}\left(l_{2}+\beta_{4}^{2}l_{3}+\beta_{4}m\right)+\beta_{4}m\left(l_{2}+\beta_{4}^{2}l_{3}+\beta_{4}m\right)\\ \end{array}\right)\begin{pmatrix}c_{2}\\ c_{4}\end{pmatrix}=0,$
where
$r_{1}=\beta_{2}^{2}l_{1}\left(l_{2}+\beta_{2}^{2}l_{3}+\beta_{2}m\right)+\beta_{2}^{2}l_{3}\left(\beta_{2}^{2}(-l_{1})-l_{1}-\beta_{2}n\right)+\beta_{2}m\left(\beta_{2}^{2}(-l_{1})-l_{1}-\beta_{2}n\right)+\beta_{2}n\left(l_{2}+\beta_{2}^{2}l_{3}+\beta_{2}m\right),$
$r_{2}=\beta_{4}^{2}l_{1}\left(l_{2}+\beta_{4}^{2}l_{3}+\beta_{4}m\right)+\beta_{4}^{2}l_{3}\left(\beta_{4}^{2}(-l_{1})-l_{1}-\beta_{4}n\right)+\beta_{4}m\left(\beta_{4}^{2}(-l_{1})-l_{1}-\beta_{4}n\right)+\beta_{4}n\left(l_{2}+\beta_{4}^{2}l_{3}+\beta_{4}m\right).$
@@
$l_{2}=l_{3}$ renders either $|\beta_{2}||\beta_{4}|=1$ or $|\beta_{2}|=|\beta_{4}|=1$, contradicting the assumption that $|\beta_{2}|<1,|\beta_{4}|<1$. Therefore, if $n(l_{2}+l_{3})=2l_{1}m$, there exists a boundary zero mode satisfying boundary condition Eq. (94). The above discussion has ignored $m^{2}=4l_{2}l_{3}$. The condition for nontrivial solution to boundary condition Eq. (94) enforces $|\beta_{2}|=|\beta_{4}|=1$, contradicting $|\beta_{2}|<1,|\beta_{4}|<1$.
The next step is the discussion on the boundary bases $\mathbf{w}^{1}$ and $\mathbf{w}^{3}$ corresponding to $\beta_{1}$ and $\beta_{3}$, respectively.
$\mathbf{w}^{1}=\begin{pmatrix}\beta_{1}w^{1,a}&\beta_{1}w^{1,b}&\beta_{1}^{2}w^{1,a}&\beta_{1}^{2}w^{1,b}&\cdots&\beta_{1}^{N}w^{1,a}&\beta_{1}^{N}w^{1,b}\end{pmatrix}^{T},$
$\mathbf{w}^{3}=\begin{pmatrix}\beta_{3}w^{3,a}&\beta_{3}w^{3,b}&\beta_{3}^{2}w^{3,a}&\beta_{3}^{2}w^{3,b}&\cdots&\beta_{3}^{N}w^{3,a}&\beta_{3}^{N}w^{3,b}\end{pmatrix}^{T}.$
@@
A trial solution $\mathbf{u}^{1,3}$ is a linear combination of $\mathbf{w}^{1}$ and $\mathbf{w}^{3}$.
@@
$\beta_{1}\neq\beta_{3}$ is discussed in the first step, which implies $m^{2}\neq 4l_{1}l_{2}$. Since $|\beta_{1}|>1$, $|\beta_{3}|>1$, $\mathbf{u}^{1,3}$ is required to satisfy the boundary condition Eq. (95), rendering
$\begin{pmatrix}\beta_{1}^{N-1}(l_{2}+\beta_{1}m)&\beta_{3}^{N-1}(l_{2}+\beta_{3}m)\\ 0&0\end{pmatrix}\begin{pmatrix}c_{1}\\ c_{3}\end{pmatrix}=0$ (149)
@@
where $c$ is a normalization constant. Therefore, there always exist boundary states satisfying the boundary condition Eq. (95).
If $m^{2}=4l_{2}l_{3}$, $\beta_{1}=\beta_{3}=\frac{-m}{2l_{3}}$. The condition for nontrivial solution to boundary condition Eq. (95) enforces $|\beta_{1}|=|\beta_{3}|=1$, contradicting $|\beta_{1}|>1,|\beta_{3}|>1$.
@@
Denote the boundary bases corresponding to $\beta_{1}$ and $\beta_{3}$ as $\mathbf{w}^{1}$ and $\mathbf{w}^{3}$, respectively.
$\mathbf{w}^{1}=\begin{pmatrix}\beta_{1}w^{1,a}&\beta_{1}w^{1,b}&\beta_{1}^{2}w^{1,a}&\beta_{1}^{2}w^{1,b}&\cdots&\beta_{1}^{N}w^{1,a}&\beta_{1}^{N}w^{1,b}\end{pmatrix}^{T},$
$\mathbf{w}^{3}=\begin{pmatrix}\beta_{3}w^{3,a}&\beta_{3}w^{3,b}&\beta_{3}^{2}w^{3,a}&\beta_{3}^{2}w^{3,b}&\cdots&\beta_{3}^{N}w^{3,a}&\beta_{3}^{N}w^{3,b}\end{pmatrix}^{T}.$
The bulk recurrence relation Eq. (86) gives
Eq. (150) is solved by
$w^{j,a}=0,\quad w^{j,b}=1,\quad j=1,3.$ (151)
A trial solution $\mathbf{u}^{1,3}$ is a linear combination of $\mathbf{w}^{1}$ and $\mathbf{w}^{3}$.
where
$\beta_{1}\neq\beta_{3}\Leftrightarrow m^{2}\neq 4l_{2}l_{3}$ is discussed at first. Since $|\beta_{1}|<1$, $|\beta_{3}|<1$, $\mathbf{u}^{1,3}$ is required to satisfy the boundary condition Eq. (94), which renders
Eq. (155) always has nontrivial solutions to $c_{1}$ and $c_{3}$.
where $c$ is a normalization constant. Therefore, there always exist boundary states satisfying the boundary condition Eq. (94). If $m^{2}=4l_{2}l_{3}$, $\beta_{1}=\beta_{3}=\frac{-m}{2l_{3}}$, the condition for nontrivial solution to boundary condition Eq. (94) enforces $|\beta_{1}|=|\beta_{3}|=1$, contradicting $|\beta_{1}|<1,|\beta_{3}|<1$.
The boundary bases corresponding to $\beta_{2}$ and $\beta_{4}$ are $\mathbf{w}^{2}$ and $\mathbf{w}^{4}$, given by
$\mathbf{w}^{2}=\begin{pmatrix}\beta_{2}w^{2,a}&\beta_{2}w^{2,b}&\beta_{2}^{2}w^{2,a}&\beta_{2}^{2}w^{2,b}&\cdots&\beta_{2}^{N}w^{2,a}&\beta_{2}^{N}w^{2,b}\end{pmatrix}^{T},$
$\mathbf{w}^{4}=\begin{pmatrix}\beta_{4}w^{4,a}&\beta_{4}w^{4,b}&\beta_{4}^{2}w^{4,a}&\beta_{4}^{2}w^{4,b}&\cdots&\beta_{4}^{N}w^{4,a}&\beta_{4}^{N}w^{4,b}\end{pmatrix}^{T}.$
@@
A trial solution $\mathbf{u}^{2,4}$ is a linear combination of $\mathbf{w}^{2}$ and $\mathbf{w}^{4}$.
@@
Since $|\beta_{2}|>1$, $|\beta_{4}|>1$, $\mathbf{u}^{2,4}$ is required to satisfy the boundary condition Eq. (95). If $\beta_{2}\neq\beta_{4}$, i.e., $m^{2}\neq 4l_{1}l_{2}$, nontrivial solution to the boundary condition Eq. (95) renders
@@
Therefore, if $n(l_{2}+l_{3})=2l_{1}m$, there exists a boundary zero mode $\mathbf{u}^{2,4}$ satisfying the boundary condition Eq. (95). $l_{2}=l_{3}$ renders $|\beta_{2}|=|\beta_{4}|=1$ or $|\beta_{2}||\beta_{4}|=1$, contradicting the assumption that $|\beta_{2}|>1,|\beta_{4}|>1$. If $m^{2}=4l_{2}l_{3}$, $\beta_{2}=\beta_{4}=\frac{-m}{2l_{2}}$, the condition for nontrivial solution to boundary condition Eq. (95) enforces $|\beta_{2}|=|\beta_{4}|=1$, contradicting $|\beta_{2}|>1,|\beta_{4}|>1$.

### v5

@@ Appendix B The Conditions for the Existence of Boundary Zero Modes under $\det B$=0
then a boundary zero mode $\mathbf{u}^{2}=\mathbf{w}^{2}$ given by Eq. (221) exists that match the boundary condition Eq. (151b), where
$w^{2,a}=\frac{l_{1}l_{3}-\frac{mn}{2}}{\beta_{2}\sqrt{\left(l_{1}l_{3}-\frac{mn}{2}\right)^{2}+\left(l_{3}^{2}-m^{2}\right)^{2}}}\sqrt{\frac{1-\beta_{2}^{2}}{1-\beta_{2}^{2N}}},\ w^{2,b}=\frac{l_{3}^{2}-m^{2}}{\beta_{2}\sqrt{\left(l_{1}l_{3}-\frac{mn}{2}\right)^{2}+\left(l_{3}^{2}-m^{2}\right)^{2}}}\sqrt{\frac{1-\beta_{2}^{2}}{1-\beta_{2}^{2N}}},\ \beta_{2}=-\frac{l_{3}}{m}.$ (260)
B.3 The conditions for boundary zero modes under $\det B=0$
The set operation $\mathbb{A}_{1}\cup\mathbb{A}_{2}\cup\mathbb{A}_{3}$ yields the conditions for the appearance of a boundary zero mode that satisfies the boundary condition Eq. (151a):
$\mathbb{A}_{1}:~{}\left\{\begin{array}[]{l}m\neq 0,\quad l_{4}\neq 0,\quad l_{2}\neq l_{3},\\ l_{1}l_{4}=l_{2}l_{3},\quad 2l_{3}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\end{array}\right.\quad\text{or}\quad\mathbb{A}_{2}:~{}\left\{\begin{array}[]{l}m\neq 0,\quad l_{3}\neq 0,\\ l_{4}=l_{2}=0,\\ |m|<|l_{3}|,\end{array}\right.\quad\text{or}\quad\mathbb{A}_{3}:~{}\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq 0,\\ nl_{2}=2l_{1}m,\quad l_{3}=l_{4}=0,\\ |m|<|l_{2}|.\end{array}\right.$ (270)
Likewise, the set operation $\mathbb{B}_{1}\cup\mathbb{B}_{2}\cup\mathbb{B}_{3}$ yields the conditions for the appearance of a boundary zero mode that satisfies the boundary condition Eq. (151b):
$\mathbb{B}_{1}:~{}\left\{\begin{array}[]{l}m\neq 0,\quad l_{4}\neq 0,\quad l_{2}\neq l_{3},\\ l_{1}l_{4}=l_{2}l_{3},\quad 2l_{2}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\end{array}\right.\quad\text{or}\quad\mathbb{B}_{2}:~{}\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq 0,\\ l_{4}=l_{3}=0,\\ |m|<|l_{2}|,\end{array}\right.\quad\text{or}\quad\mathbb{B}_{3}:~{}\left\{\begin{array}[]{l}m\neq 0,\quad l_{3}\neq 0,\\ nl_{3}=2l_{1}m,\quad l_{2}=l_{4}=0,\\ |m|<|l_{3}|.\end{array}\right.$ (280)
Taking all the conditions for boundary zero modes into account, we conclude that the $K$ matrix in Eq. (2) has one boundary zero mode if one of the following conditions is satisfied:
$\left\{\begin{array}[]{l}m\neq 0,\quad l_{4}\neq 0,\\ l_{1}l_{4}=l_{2}l_{3},\\ 2l_{3}m=l_{4}n\ \text{ or }\ 2l_{2}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\end{array}\right.\text{ or }\left\{\begin{array}[]{l}m\neq 0,\quad l_{3}\neq 0,\\ l_{4}=l_{2}=0,\\ nl_{3}\neq 2l_{1}m\\ |m|<|l_{3}|,\end{array}\right.\text{ or }\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq 0,\\ l_{4}=l_{3}=0,\\ nl_{2}\neq 2l_{1}m\\ |m|<|l_{2}|,\end{array}\right.$ (293)
@@ B.3 The conditions for boundary zero modes under $\det B=0$
which is obtained by the set operation $\mathbb{V}\backslash[(\mathbb{A}_{1}\cup\mathbb{A}_{2}\cup\mathbb{A}_{3})\cap(\mathbb{B}_{1}\cup\mathbb{B}_{2}\cup\mathbb{B}_{3})]$, where $\mathbb{V}=\mathbb{A}_{1}\cup\mathbb{A}_{2}\cup\mathbb{A}_{3}\cup\mathbb{B}_{1}\cup\mathbb{B}_{2}\cup\mathbb{B}_{3}$. More specifically, if $\left\{\begin{array}[]{l}l_{4}=0\\ |l_{2}|<|l_{3}|\end{array}\right.\text{ or }\left\{\begin{array}[]{l}l_{4}\neq 0\\ 2l_{3}m=l_{4}n\end{array}\right.$, the boundary zero mode is a lower boundary zero mode, satisfying the boundary condition Eq. (151a). If $\left\{\begin{array}[]{l}l_{4}=0\\ |l_{3}|<|l_{2}|\end{array}\right.\text{ or }\left\{\begin{array}[]{l}l_{4}\neq 0\\ 2l_{2}m=l_{4}n\end{array}\right.$, the boundary zero mode is an upper boundary zero mode, satisfying the boundary condition Eq. (151b).
Moreover, if the elements of the $K$ matrix satisfies
$\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq 0,\\ l_{3}=l_{4}=0,\\ nl_{2}=2l_{1}m,\\ |m|<|l_{2}|,\end{array}\right.\text{ or }\left\{\begin{array}[]{l}m\neq 0,\quad l_{3}\neq 0,\\ l_{2}=l_{4}=0,\\ nl_{3}=2l_{1}m,\\ |m|<|l_{3}|,\end{array}\right.$ (302)
@@
the $K$ matrix has two boundary zero modes, which is obtained by the set operation $(\mathbb{A}_{1}\cup\mathbb{A}_{2}\cup\mathbb{A}_{3})\cap(\mathbb{B}_{1}\cup\mathbb{B}_{2}\cup\mathbb{B}_{3})$.
Appendix C The Conditions for the Existence of Boundary Zero Modes under $\det B\neq 0$
The parameters $\{\beta_{j}\}$ in the boundary base (Eq. (143)) controls the boundary condition that a trial solution (Eq. (142)) is required to satisfy. In the discussion of $\det B\neq 0$ cases, the solutions for $\beta_{j}$ ($j=1,2,3,4$) are provided by Eq. (146), which is a quartic equation and the solutions for $\beta$ are given by:
@@ Appendix C The Conditions for the Existence of Boundary Zero Modes under $\det B\neq 0$
These solutions render the relation
@@
If $l_{4}\neq 0$, carrying on a comprehensive analytical analysis is challenging and requires a case-by-case study, which is left for future research. To simplify discussion, we focus on the special case where $l_{4}=0$. Setting $l_{4}=0$, Eq. (146) becomes
$(l_{2}+m\beta+l_{3}\beta^{2})(l_{3}+m\beta+l_{2}\beta^{2})=0.$ (304)
$\det B\neq 0$ together with $l_{4}=0$ renders $l_{2}l_{3}\neq 0$. Hence the solutions to Eq. (304) are
@@
Case I. $|\beta_{1}|>1$, $|\beta_{2}|<1$, $|\beta_{3}|>1$, $|\beta_{4}|<1$;
Case II. $|\beta_{1}|<1$, $|\beta_{2}|>1$, $|\beta_{3}|<1$, $|\beta_{4}|>1$;
Case III. $|\beta_{1}|>1$, $|\beta_{2}|<1$, $|\beta_{3}|<1$, $|\beta_{4}|>1$;
@@
In the subsequent analysis, we divide our discussion into two subsections addressing boundary conditions Eq. (151a) and Eq. (151b), respectively. Within each subsection, the discussion is organized into four distinct cases: Case I $\sim$ Case IV. For each case, trial solutions are constructed using the boundary bases. Each trial solution, denoted as $\mathbf{u}^{i,j}$, is a linear superposition of two boundary bases, $\mathbf{w}^{i}$ and $\mathbf{w}^{j}$ (Eq. (143)), where the moduli of parameters $\beta_{i}$ and $\beta_{j}$ are either both greater than 1 or both less than 1. These trial solutions are required to match the corresponding boundary conditions. This procedure yields explicit constraints on the elements of the $K$ matrix that guarantee the existence of boundary zero modes.
C.1 Discussion on the boundary condition Eq. (151a)
@@
Since $|\beta_{2}|<1$, $|\beta_{4}|<1$, $\mathbf{u}^{2,4}$ is required to satisfy the boundary condition Eq. (151a). The boundary bases corresponding to $\beta_{2}$ and $\beta_{4}$ are denoted as $\mathbf{w}^{2}$ and $\mathbf{w}^{4}$, respectively.
$\mathbf{w}^{j}=\begin{pmatrix}\beta_{j}w^{j,a}&\beta_{j}w^{j,b}&\beta_{j}^{2}w^{j,a}&\beta_{j}^{2}w^{j,b}&\cdots&\beta_{j}^{N}w^{j,a}&\beta_{j}^{N}w^{j,b}\end{pmatrix}^{T},\quad j=2,4.$
@@
$\mathbf{u}^{2,4}$ is a linear combination of $\mathbf{w}^{2}$ and $\mathbf{w}^{4}$.
where
@@
The boundary condition Eq. (151a) renders
$\left(\begin{array}[]{cc}\beta_{2}^{2}\left(l_{1}l_{2}-l_{1}l_{3}\right)+\beta_{2}\left(l_{2}n-l_{1}m\right)&\beta_{4}^{2}\left(l_{1}l_{2}-l_{1}l_{3}\right)+\beta_{4}\left(l_{2}n-l_{1}m\right)\\ \beta_{2}^{2}l_{2}\left(l_{2}+\beta_{2}^{2}l_{3}+\beta_{2}m\right)+\beta_{2}m\left(l_{2}+\beta_{2}^{2}l_{3}+\beta_{2}m\right)&\beta_{4}^{2}l_{2}\left(l_{2}+\beta_{4}^{2}l_{3}+\beta_{4}m\right)+\beta_{4}m\left(l_{2}+\beta_{4}^{2}l_{3}+\beta_{4}m\right)\\ \end{array}\right)\begin{pmatrix}c_{2}\\ c_{4}\end{pmatrix}=0.$
@@
$l_{2}=l_{3}$ renders either $|\beta_{2}||\beta_{4}|=1$ or $|\beta_{2}|=|\beta_{4}|=1$, contradicting the assumption that $|\beta_{2}|<1,|\beta_{4}|<1$. Therefore, if the elements of the $K$ matrix satisfy
$\mathbb{C}_{1}:~{}\left\{\begin{array}[]{l}l_{4}=0,\\ m\neq 0,\quad l_{2}\neq l_{3},\quad l_{2}l_{3}\neq 0\\ |\beta_{1}|>1,|\beta_{2}|<1,|\beta_{3}|>1,|\beta_{4}|<1,\\ n(l_{2}+l_{3})=2l_{1}m,\end{array}\right.$ (311)
there exists a boundary zero mode satisfying boundary condition Eq. (151a).
Case II
$|\beta_{1}|<1$, $|\beta_{2}|>1$, $|\beta_{3}|<1$, $|\beta_{4}|>1$.
Since $|\beta_{1}|<1$, $|\beta_{3}|<1$, $\mathbf{u}^{1,3}$ is required to satisfy the boundary condition Eq. (151a). Denote the boundary bases corresponding to $\beta_{1}$ and $\beta_{3}$ as $\mathbf{w}^{1}$ and $\mathbf{w}^{3}$, respectively.
$\mathbf{w}^{j}=\begin{pmatrix}\beta_{j}w^{j,a}&\beta_{j}w^{j,b}&\beta_{j}^{2}w^{j,a}&\beta_{j}^{2}w^{j,b}&\cdots&\beta_{j}^{N}w^{j,a}&\beta_{j}^{N}w^{j,b}\end{pmatrix}^{T},\quad j=1,3.$
The bulk recurrence relation Eq. (144) gives
Eq. (312) is solved by
$w^{j,a}=0,\quad w^{j,b}=1,\quad j=1,3.$ (313)
$\mathbf{u}^{1,3}$ is a linear combination of $\mathbf{w}^{1}$ and $\mathbf{w}^{3}$.
where
The boundary condition Eq. (151a) renders
Eq. (317) always has nontrivial solutions to $c_{1}$ and $c_{3}$.
where $c$ is a normalization constant. Therefore, there exists a boundary zero mode satisfying the boundary condition Eq. (151a) under the condition
$\mathbb{C}_{2}:~{}\left\{\begin{array}[]{l}l_{4}=0,\\ m\neq 0,\quad l_{2}l_{3}\neq 0,\\ |\beta_{1}|<1,|\beta_{2}|>1,|\beta_{3}|<1,|\beta_{4}|>1.\end{array}\right.$ (322)
Case III
$|\beta_{1}|>1$, $|\beta_{2}|<1$, $|\beta_{3}|<1$, $|\beta_{4}|>1$.
Since $|\beta_{2}|<1,|\beta_{3}|<1$, $\mathbf{u}^{2,3}$ is required to satisfy the boundary condition specified by Eq. (151a). The boundary bases corresponding to $\beta_{2}$ and $\beta_{3}$ are denoted as $\mathbf{w}^{2}$ and $\mathbf{w}^{3}$, respectively.
$\mathbf{w}^{j}=\begin{pmatrix}\beta_{j}w^{j,a}&\beta_{j}w^{j,b}&\beta_{j}^{2}w^{j,a}&\beta_{j}^{2}w^{j,b}&\cdots&\beta_{j}^{N}w^{j,a}&\beta_{j}^{N}w^{j,b}\end{pmatrix}^{T},\quad j=2,3.$
The bulk recurrence relation Eq. (144) gives
The solution to Eq. (323) is given by
$\mathbf{u}^{2,3}$ is a linear combination of $\mathbf{w}^{2}$ and $\mathbf{w}^{3}$.
where
The boundary condition Eq. (151a) renders
Nontrivial solution to Eq. (326) implies
Eq. (329) together with Eq. (305b), Eq. (305c) renders
Eq. (330) together with Eq. (305b) and Eq. (305c) enforces $|\beta_{2}|=1$, which contradicts our assumption $|\beta_{2}|<1,|\beta_{3}|<1$.
Case IV
$|\beta_{1}|<1$, $|\beta_{2}|>1$, $|\beta_{3}|>1$, $|\beta_{4}|<1$.
Since $|\beta_{1}|<1,|\beta_{4}|<1$, $\mathbf{u}^{1,4}$ is required to satisfy the boundary condition specified by Eq. (151a). Denote the boundary bases corresponding to $\beta_{1}$ and $\beta_{4}$ as $\mathbf{w}^{1}$ and $\mathbf{w}^{4}$, respectively.
$\mathbf{w}^{j}=\begin{pmatrix}\beta_{j}w^{j,a}&\beta_{j}w^{j,b}&\beta_{j}^{2}w^{j,a}&\beta_{j}^{2}w^{j,b}&\cdots&\beta_{j}^{N}w^{j,a}&\beta_{j}^{N}w^{j,b}\end{pmatrix}^{T},\quad j=1,4.$
The bulk recurrence relation Eq. (144) gives
Eq. (331) is solved by
$\mathbf{u}^{1,4}$ is a linear combination of $\mathbf{w}^{1}$ and $\mathbf{w}^{4}$.
where
The boundary condition Eq. (151a) renders
Nontrivial solution to Eq. (334) requires
Eq. (337) along with Eq. (305a) and Eq. (305d) renders
Eq. (338), Eq. (305a) and Eq. (305d) enforce $|\beta_{4}|=1$, which contradicts our assumption $|\beta_{1}|<1$, $|\beta_{4}|<1$.
C.2 Discussion on the boundary condition Eq. (151b)
Case I
$|\beta_{1}|>1$, $|\beta_{2}|<1$, $|\beta_{3}|>1$, $|\beta_{4}|<1$.
Since $|\beta_{1}|>1$, $|\beta_{3}|>1$, $\mathbf{u}^{1,3}$ is required to satisfy the boundary condition Eq. (151b). The boundary bases corresponding to $\beta_{1}$ and $\beta_{3}$ are $\mathbf{w}^{1}$ and $\mathbf{w}^{3}$, respectively.
$\mathbf{w}^{j}=\begin{pmatrix}\beta_{j}w^{j,a}&\beta_{j}w^{j,b}&\beta_{j}^{2}w^{j,a}&\beta_{j}^{2}w^{j,b}&\cdots&\beta_{j}^{N}w^{j,a}&\beta_{j}^{N}w^{j,b}\end{pmatrix}^{T},\quad j=1,3.$
@@
$\mathbf{u}^{1,3}$ is a linear combination of $\mathbf{w}^{1}$ and $\mathbf{w}^{3}$.
@@
The boundary condition Eq. (151b) yields
$\begin{pmatrix}\beta_{1}^{N-1}(l_{2}+\beta_{1}m)&\beta_{3}^{N-1}(l_{2}+\beta_{3}m)\\ 0&0\end{pmatrix}\begin{pmatrix}c_{1}\\ c_{3}\end{pmatrix}=0.$ (340)
@@
where $c$ is a normalization constant. Therefore, under the condition
$\mathbb{D}_{1}:~{}\left\{\begin{array}[]{l}l_{4}=0,\\ m\neq 0,\quad l_{2}l_{3}\neq 0,\\ |\beta_{1}|>1,|\beta_{2}|<1,|\beta_{3}|>1,|\beta_{4}|<1,\end{array}\right.$ (344)
there always exists a boundary zero mode satisfying the boundary condition Eq. (151b).
@@
Since $|\beta_{2}|>1$, $|\beta_{4}|>1$, $\mathbf{u}^{2,4}$ is required to satisfy the boundary condition Eq. (151b). The boundary bases corresponding to $\beta_{2}$ and $\beta_{4}$ are $\mathbf{w}^{2}$ and $\mathbf{w}^{4}$, given by
$\mathbf{w}^{j}=\begin{pmatrix}\beta_{j}w^{j,a}&\beta_{j}w^{j,b}&\beta_{j}^{2}w^{j,a}&\beta_{j}^{2}w^{j,b}&\cdots&\beta_{j}^{N}w^{j,a}&\beta_{j}^{N}w^{j,b}\end{pmatrix}^{T},\quad j=2,4.$
@@
$\mathbf{u}^{2,4}$ is a linear combination of $\mathbf{w}^{2}$ and $\mathbf{w}^{4}$.
@@
The boundary condition Eq. (151b) yields
$\left(\begin{array}[]{cc}-\beta_{2}^{N+1}\left(l_{1}\left(l_{2}-l_{3}+\beta_{2}m\right)-\beta_{2}l_{3}n\right)&-\beta_{4}^{N+1}\left(l_{1}\left(l_{2}-l_{3}+\beta_{4}m\right)-\beta_{4}l_{3}n\right)\\ \beta_{2}^{N-1}\left(l_{3}+\beta_{2}m\right)\left(\beta_{2}\left(\beta_{2}l_{3}+m\right)+l_{2}\right)&\beta_{4}^{N-1}\left(l_{3}+\beta_{4}m\right)\left(\beta_{4}\left(\beta_{4}l_{3}+m\right)+l_{2}\right)\\ \end{array}\right)\begin{pmatrix}c_{2}\\ c_{4}\end{pmatrix}=0.$ (347)
Nontrivial solution to the boundary condition Eq. (151b) renders
@@
$l_{2}=l_{3}$ renders $|\beta_{2}|=|\beta_{4}|=1$ or $|\beta_{2}||\beta_{4}|=1$, which contradicts the assumption that $|\beta_{2}|>1,|\beta_{4}|>1$. Therefore, if
$\mathbb{D}_{2}:~{}\left\{\begin{array}[]{l}l_{4}=0,\\ m\neq 0,\quad l_{2}\neq l_{3},\quad l_{2}l_{3}\neq 0\\ |\beta_{1}|<1,|\beta_{2}|>1,|\beta_{3}|<1,|\beta_{4}|>1,\\ n(l_{2}+l_{3})=2l_{1}m,\end{array}\right.$ (352)
there exists a boundary zero mode $\mathbf{u}^{2,4}$ satisfying the boundary condition Eq. (151b).